Please help me solve this problem: Maria is 18 years old. She is twice as old as Anna was when she was as old as Anna now. How old is Anna now?
Answer by ptaylor(2198)

Maria is 18 years old now
Let x=Anna's age now
Let y=number of years that have passed since Maria was x years old
Sooo
18-y=x
x+y=18---eq1
Then when Maria was x years old , Anna was x-y years old
Sooo
18=2(x-y)=2x-2y or
x-y=9----eq2
Add eq1 and eq2 and we get:
2x=27
x=13.5 Anna's age now
substitute 13.5 into eq1
y=18-13.5=4.5 number of years that have passed since Maria was 13.5 years old. So, when Maria was 13.5, Anna was 13.5-4.5=9
18=2*9
Hope this helps---ptaylor